OpenStudy (anonymous):

Solve this inequality. -9x - 3 > or = to 51

OpenStudy (anonymous):

\[-9x - 3 \ge 51\] first you have to put all the known values on one side of the equation and all variables on the other side \[-9x \ge 51 + 3\] when we transfer 3 from left to right, minus in front changes to plus \[-9x \ge 54\] now you just have to divide whole inequalities with -9 (also, remember that when you divide inequalities with negative number, sign of inequality changes to the opposite, i.e. > or = becomes <)

OpenStudy (anonymous):

ok thanks